OpenStudy (anonymous):

If y – 1 = 4x, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? plz help will give medal {(1, 4), (2, 8), (3, 12)} {(4, 1), (8, 2), (12, 3)} {(0, 1), (1, 5), (2, 9)} {(1, 0), (5, 1), (9, 2)}

OpenStudy (dtan5457):

y-1=4x add one y=4x+1 now just plug each x coordinate of each answer and see if the y input is correct.

OpenStudy (anonymous):

would i add 1 to 2,8 to ??

OpenStudy (anonymous):

or do i add the 8

OpenStudy (dtan5457):

This is your function y=4x+1 if i wanted to test the coordinate (1,4) I'd plug 1 as x and see if 4 becomes y. y=4(1)+1 y=5 1,4 is a invalid solution

OpenStudy (anonymous):

oh i get it ok thx